Reports 1-1 of 1 Clear search Modify search
AdV-PAY (Optical Levers)
ruggi - 18:37 Wednesday 19 January 2022 (54534) Print this report
Jump in NI marionette optical lever during TCS activity

This morning we had some trouble in the recovery of north arm alignment. The reason was a large misalignment of NI, in a single step occurred at about 10:30 UTC (fig 1). The interpretation is that a fake offset was added to MAR signals, then the controls moved the payload in a different position, as seen by MIR optical lever (out of loop). The event is visible also in the optical lever power; it is very fast and it is coincident with some TCS CO2 turning on (fig 2).

In principle there is nothing warrying in the fact that we have a new offset in the optical lever, but after the event a line at around 5 Hz appeared in the signals. We are not aware of a mechanical resonance at that frequency. Moreover, the line is persistent, has several large harmonics and is not stable in frequency, therefore it could be due to a disturbance entering the sensor. It is worth to identify the source of the step and the disturbance.

Images attached to this report
Comments to this report:
fafone, rocchi - 23:14 Wednesday 19 January 2022 (54540) Print this report

The switch on of the NI CO2 laser seems to have affected also other signals around in a non-predictable manner, too fast to be thermal. To be checked.

Images attached to this comment
Paoletti, Fiori, Tringali - 15:23 Thursday 20 January 2022 (54547) Print this report

We are lucky enough to have three magnetometers all around NI base tower. If there is any magnetic signal appearing @ 5Hz, they should be seeing it for sure.

Indeed there is. It appeared at 5.2 Hz, decaying to around 4.3 Hz after few hours. The sensor that see it better is ENV_CEB_MAG_W


Images attached to this comment
Paoletti - 12:19 Friday 21 January 2022 (54561) Print this report

A better witness to this new noise appears to be ENV_CEB_ELECTRIC  (the channel that monitors the voltage drop between the Virgo vacuum steel tube and the mains Protection Earth system).

The signal appeared at 10:31:15 UTC and this should coincide with the power-ON of the 40W NI CO2 laser (TBC)

Power-ON occurs with two actions: apply the 40 MHz RF signal from the RF generator to the RF amplifier and apply the 5V enable voltage to the laser head.

These two actions were done almost simultaneously, they should be repeated to figure out who is responsible for the new noise.

Images attached to this comment
cavalieri, dattilo, passaquieti - 11:15 Wednesday 26 January 2022 (54609) Print this report

After the replacement of the RF driver, the switch on of the NI CO2 laser affected also signals acquired by IMMS and ENV systems.

In particular (see jump3.gif) the jump is visible in INF_TCS_CHILROOM_TE (AD590 temperature probe inside TCS chiller room).

Also temperature sensors (AD590) installed into NI superattenuators see a similar jump.

In jump4.gif we see the effect of switch on of the NI CO2 after the replacement of the RF driver with the original one.

The INF_TCS_CHILROOM_TE signal doesn't  change because in the meanwhile we replaced the sensor with a PT100, while NI superattenuator temperature sensor are still affected by the jump.


Images attached to this comment
dattilo - 12:38 Wednesday 26 January 2022 (54610) Print this report
In the various plots of the temperature channels, each jump is characterized by a transient with overshot: this is an artifact due to the effect of the anti-aliasing filter (8th order, cut-off frequency = 0.1Hz) in presence of the sudden jump in the signals.
Search Help